OpenStudy (anonymous):

Help please f(x)=2-sqrtx/2+3 What are my x and y intercepts and domain and range? I think I may have calculated something wrong :(

OpenStudy (xapproachesinfinity):

y intercept st x=0 x intercept set y=0 mean f(x)=0 and solve the equation domain is the function you wrote like this \(\Large \bf\color{blueviolet}{f(x)=2-\sqrt{\frac{x}{2}+3}}\)

OpenStudy (xapproachesinfinity):

@Kerriy check the function not sure what you wrote

OpenStudy (anonymous):

Yes that is the function :)

OpenStudy (xapproachesinfinity):

Okay for domain is any number x such that \(\Large \sf\color{blueviolet}{\frac{x}{2}+3\geq0}\) solve this inequality to get the domain

OpenStudy (anonymous):

\[x \ge-6\]?

OpenStudy (jdoe0001):

\(\large { f(x)=2-\sqrt{\frac{x}{2}+3} \\ \quad \\ \quad \begin{cases} y-intercept\to y=2-\sqrt{\frac{{\color{brown}{ 0}}}{2}+3}\to y=2-\sqrt{3} \\ \quad \\ x-intercept\to {\color{brown}{ 0}}=2-\sqrt{\frac{x}{2}+3}\to \sqrt{\frac{x}{2}+3}=2 \\ \quad \\ \frac{x}{2}+3=2^2\to \frac{x}{2}=4-3\to x=2(4-3) \end{cases} }\)

OpenStudy (xapproachesinfinity):

y-intercept is easy just evaluate 0 \(\Large \sf\color{blueviolet}{f(0)=2-\sqrt{\frac{0}{2}+3}=2-\sqrt{3}}\)

OpenStudy (xapproachesinfinity):

Correct that's for domain meaning all real numbers that greater than or equal to -6

OpenStudy (xapproachesinfinity):

You got the answer for everything!

OpenStudy (anonymous):

what would my starting point be?

OpenStudy (xapproachesinfinity):

what do you mean by starting point?

OpenStudy (anonymous):

no sorry my end point

OpenStudy (xapproachesinfinity):

end point for what, i don't get what you are trying to ask

OpenStudy (anonymous):

Doesn't matter, I think I have worked it out.....thanks so very much for all your help :)

OpenStudy (xapproachesinfinity):

\(\Large \frak\color{blueviolet}{you~welcome}\)